# PLS-complete problems with lexicographic cost functions: Max-$k$-SAT and Abelian Permutation Orbit Minimization

### v1

@@ Abstract
How hard is it to find a local optimum? If we are given a graph and want to find a locally maximal cut–meaning that the number of edges in the cut cannot be improved by moving a single vertex from one side to the other–then just iterating improving steps finds a local maximum since the size of the cut can increase at most $|E|$ times. If, on the other hand, the edges are weighted, this problem becomes hard for the class PLS (Polynomial Local Search) [15].
@@
We use these results to answer a question in Scheder and Tantow [16], who showed that finding a lexicographic local minimum of a string $s\in\{0,1\}^{n}$ under the action of a list of given permutations $\pi_{1},\dots,\pi_{k}\in S_{n}$ is PLS-complete. They ask whether the problem stays PLS-complete when the $\pi_{1},\dots,\pi_{k}$ commute, i.e., generate an Abelian subgroup $G$ of $S_{n}$. In this work, we show that it does, and in fact stays PLS-complete even (1) when every element in $G$ has order two and also (2) when $G$ is cyclic, i.e., all $\pi_{1},\dots,\pi_{k}$ are powers of a single permutation $\pi$.
@@ 1 Introduction
is the lexicographically smallest? This problem is known to be NP-hard even for $m=1$ [16]. A more modest goal is to find a local minimum, i.e., a $v=s\circ\sigma$ for some $\sigma\in\langle\pi_{1},\dots,\pi_{m}\rangle$ such that none of $v\circ\pi_{1},v\circ\pi_{2},\dots,v\circ\pi_{m}$ is lexicographically smaller than $v$. Such a local minimum obviously exists and can be verified efficiently (this is less obvious and we discuss it later). This means finding a local minimum is in the complexity class PLS (polynomial local search), more formally defined in the next sections. We call this problem Permutation Orbit Local Minimization, short Polm. Scheder and Tantow [16] showed that Polm is PLS-complete and thus unlikely to admit an efficient algorithm. They left it as an open question whether Polm is still PLS-complete if all given permutations commute, i.e., if $\langle\pi_{1},\dots,\pi_{m}\rangle$ is Abelian. We answer this question: Yes, Abelian Polm is still PLS-complete.
@@ 1.1 Motivation
The motivation behind Polm comes from SAT solving. Kołodziejczyk and Thapen [8] investigate whether proof systems including symmetry breaking rules can be simulated by proof systems without them. For a concrete example, suppose we want to prove that the Ramsey number $R(5,5)$ is at least $43$. The statement “$G$ is a graph on $43$ vertices and has no clique nor independent set of size 55” can be written as a CNF formula (here: a 10-CNF formula with $\binom{43}{2}$ variables) and fed to a SAT solver. If there is a solution, then every graph isomorphic to it is also a solution. Thus, we can try to speed things up by adding certain symmetry breaking constraints. For example, we might want to add constraints formalizing the statement “and $G$ should be lexicographically minimal”. If by “minimal” we mean “a global minimum” then this statement cannot be written efficiently as a CNF formula; if we mean “the graph cannot be made smaller by exchanging the names of two vertices” then we can encode this as a CNF formula; we are now in fact looking for a solution that is a lexicographic local minimum with respect to the permutations on the set of edges defined by swapping two vertices.
@@ 1.2 Total search problems
Another example of an FNP-problem would be Factoring, where $x$ is a number and $y$ is a prime factor of $x$, which can be checked in polynomial time. In contrast to SAT, every instance $x$ has at least one solution $y$ and hence it is a total search problem. These class of total search problems is called TFNP. If a TFNP-problem were to be NP-complete, then ${\textsc{NP}}={\textsc{coNP}}$ would follow[12]. Hence NP-hardness is not the right tool to study the hardness of total search problems. As TFNP is a semantic class, no complete problems are known. Instead, it is usually studied via its subclasses based on the combinatorial principle used to prove the totality of the problem. Here are some examples of subclasses, together with their defining combinatorial principle.
  • PLS: Every DAG has a sink [7].
  • PPAD: If a directed graph has a vertex $u$ with more edges going out than coming in, then there must be some other vertex $v$ with more edges coming in than going out [13]
  • PPA: If an undirected graph has a vertex $u$ of odd degree, it must have another such vertex $v$ [13].
  • PPP: The pigeonhole principle ([13]).
@@ 1.3 PLS
Local search is a powerful technique to solve optimization problems in practice. Nonetheless, there are instances where local search takes a long time to converge. In order to study local search from a theoretical perspective, the class PLS was introduced in 1988 [7]. Following [5], a combinatorial optimization problem can be stated as follows:
@@
PLS is now defined as the class of optimization problems where start solution, cost, and neighborhood are computable in polynomial time, and the set SSS of feasible solutions can be recognized efficiently. A local optimum is a feasible solution $s$ such that $\textnormal{cost}(s)\geq\textnormal{cost}(s^{\prime})$ for all $s^{\prime}\in N(s)$. [^2] Alternatively, we can also view it as a minimization problem, where the cost of $s$ is smaller or equal compared to all of its neighbors.
A local optimum can be found with the standard algorithm: We compute a starting solution in polynomial time and then always select an improving neighbor until no such neighbor exists. There are instances where this can take exponentially many steps[7]. In fact, the standard algorithm can take exponentially long even when the optimization problem itself is in P: think about linear programming and the simplex algorithm.
PLS contains problems that are complete under search problem reductions, for example finding a locally optimal max-cut in a weighted graph[15], finding locally optimal assignments for weighted CNF-formulas[9] or finding pure Nash equilibria in congestion games[3].
@@ 1.4 Lexicographic costs
Therefore, the costs in PLS-reductions must be large, usually of the form $2^{i}$. However, often multiple entries of the cost vector are the same. In order to compare two solutions, one would have to compute the costs of both and then compare them, which requires multiple additions and multiplications. If the cost function is like $c_{i}=2^{-i}$ and $M=2$, then we get a lexicographic cost function: $\textnormal{cost}(p_{1},x_{2})<\textnormal{cost}(p_{2},x_{2})$ if and only if $p_{1}<_{\rm lex}p_{2}$. Thus, the statement that $s=(p,x)$ is lexicographically minimal can be encoded by a simple CNF formula, which is important in contexts such as [8]. [^3] For larger values of $M$ we would define lexicographic costs as $c_{i}=M^{-i}$, but we do not encounter this case in our reductions. Thus, we call a cost function lexicographic if the coefficients of the cost vector fulfill $c_{i}=C\cdot M^{-i}$, where $M$ is usually $2$, and the common factor $C$ is just for convenience, to allow for a vector $(16,8,4,2,1)$, for example. If a cost function is lexicographic, then to compare $(p,x)$ and $(p^{\prime},x^{\prime})$ it suffices to look for the first position $i$ with $p_{i}\neq p^{\prime}_{i}$; everything to the right pales in comparison. Arguably, this is one of the most simple cost function available. The central question is: Which problems are PLS-complete problems when we insist on lexicographic costs?
The first known PLS-complete problem FLIP[7] fulfills this property. We are given a circuit $C$ with $n$ inputs and $m$ outputs and want to compute a locally minimal output. Formally, this optimization problem is defined by the relation $S=\{(C(x),x)\ |\ x\in\{0,1\}^{n}\}$, with the output $C(x)$ being the cost part and input $x$ being the non-cost part. The neighborhood of $(C(x),x)$ is the set of all $(C(x\oplus\mathbf{e}_{i}),x\oplus\mathbf{e}_{i})$, i.e., we are allowed to flip one bit of the input. The cost function is $c_{i}=2^{-i}$, i.e., we want to lexicographically minimize the output. To the best of our knowledge, this has for the longest time been the only known PLS-complete problem with lexicographic costs. Recently, Polm joined this club [16].
@@
Multiple PLS-complete problems are easy with a lexicographic cost function. Consider the weighted independent set problem[15]. Our greedy algorithm sorts the vertices by weight from high to low and starts with $I=$. For $i=1,\dots,n$ it adds $v_{i}$ to $I$ if $I\cup\{v_{i}\}$ is still independent. Similarly, we can find a global optimum of a max-cut instance with lexicographic weights. Here we have to check whether a given set of edges can be realized in a cut. This is easy because we just have to check whether they form a bipartite graph. More generally, Max-2-SAT with lexicographic costs can be solved efficiently because a given subset of clauses can be efficiently tested for satisfiability. It is not obvious how this works for larger clauses, as there is no known way to check satisfiability of a subset of clauses. In fact, we show that finding a local maximum under lexicographic clause weights is PLS-hard for 4-CNF formulas. We should note that even though a global lex-maximum for 2-SAT can be found efficiently, we do not know whether the standard algorithm terminates after a polynomial number of steps.
@@ 1.5 Reductions in PLS
In order to talk about PLS-completeness, we need a notion of reductions. A PLS reduction[7] from $P$ to $Q$ is a pair of functions $f$ and $g$. $f$ maps an instance $I$ of $P$ to an instance $f(I)$ of $Q$. $g$ maps a solution $s$ of $f(I)$ to a solution $g(I,s)$ of $I$, so that if $s$ is a local optimum for $Q$ then $g(I,s)$ is a local optimum for $P$. Additionally, both $f$ and $g$ are computable in polynomial time. One checks that this fulfills the purpose of a reduction: if we can efficiently solve $Q$ (i.e., find local minimum) then we can solve $P$. The transition graph $TG(I)$ of an instance $I$ of a PLS-problem $P$ is the directed graph where the vertices are the feasible solutions of $I$ and an edge $(u,v)$ is present if $v$ is a neighbor of $u$ according to the neighborhood function.
@@
A PLS-reduction $(f,g)$ from $P$ to $Q$ is called tight[15] if for every instance $I$ of $P$ there exists a set $\mathcal{R}$ of feasible solutions of the instance $f(I)$ of $Q$ such that
@@ 1.6 Previous results
The Polm problem arose from proof-complexity based investigations into the dominance rule by Kołodziejczyk and Thapen [8]. Its PLS-completeness was shown by Scheder and Tantow in [16]. If the output of Polm is the locally minimized string $x\in\{0,1\}^{n}$ itself, then the problem of verifying that $x$ is in the orbit, i.e., that there is a permutation $\pi\in\langle\pi_{1},\dots,\pi_{k}\rangle$ with $x=s\circ\pi$ is known as the string isomorphism problem, which is at least as hard as graph isomorphism. Thus, to make sure Polm is in PLS, we want the output to not be $x$ but the permutation $\pi$ itself; now whether $\pi\in\langle\pi_{1},\dots,\pi_{k}\rangle$ can be verified efficiently using an algorithm of Furst, Hopcroft, and Luks [4]. For the hardness proof, they used a direct reduction from FLIP using gate gadgets and the test circuit technique by Krentel[9]. While the gate gadgets could be built with commuting permutations, the test circuit technique introduced permutations that do not commute with the permutations of the gate gadgets. Hence, the question of the complexity of Polm for Abelian permutation groups arose.
Given a CNF formula with clause weights and an assignment $\alpha$, the payoff of $\alpha$ is the total weight of satisfied clauses. We denote by $k$-SAT/$d$-FLIP the optimization problem, where we are given a $k$-CNF formula with clause weights and want to find a truth assignment $\alpha$ whose payoff is locally optimal: switching $\alpha$ at at most $d$ variables does not increase the payoff. For arbitrary costs (i.e., not necessarily lexicographic) it has been known that $k$-SAT/1-FLIP is PLS-complete [10]. [^4] In fact, Krentel gives a reduction to $(3,4)$-CSP and then claims that it is easy to see that it also works for SAT. Later, [15] showed that Max-Cut is PLS-compete, which immediately implies that $2$-SAT/1-FLIP is PLS-complete.
Further general PLS reductions can be found in [7, 15, 9, 2]. Buchheim and Jünger proved NP-completeness for multiple problems regarding permutations in [1].

### v2

@@ Abstract
How hard is it to find a local optimum? If we are given a graph and want to find a locally maximal cut–meaning that the number of edges in the cut cannot be improved by moving a single vertex from one side to the other–then just iterating improving steps finds a local maximum since the size of the cut can increase at most $|E|$ times. If, on the other hand, the edges are weighted, this problem becomes hard for the class PLS (Polynomial Local Search) [DBLP:journals/siamcomp/SchafferY91].
@@
We use these results to answer a question in Scheder and Tantow [SchederTantow2025], who showed that finding a lexicographic local minimum of a string $s\in\{0,1\}^{n}$ under the action of a list of given permutations $\pi_{1},\dots,\pi_{k}\in S_{n}$ is PLS-complete. They ask whether the problem stays PLS-complete when the $\pi_{1},\dots,\pi_{k}$ commute, i.e., generate an Abelian subgroup $G$ of $S_{n}$. In this work, we show that it does, and in fact stays PLS-complete even (1) when every element in $G$ has order two and also (2) when $G$ is cyclic, i.e., all $\pi_{1},\dots,\pi_{k}$ are powers of a single permutation $\pi$.
Additionally, we use it to reprove the hardness of computing an $\alpha$ approximate Nash equilibria in congestion games by Skopalik and Vöcking[DBLP:conf/stoc/SkopalikV08] and extend the result from step functions to exponential functions.
@@ 1 Introduction
is the lexicographically smallest? This problem is known to be NP-hard even for $m=1$ [SchederTantow2025]. A more modest goal is to find a local minimum, i.e., a $v=s\circ\sigma$ for some $\sigma\in\langle\pi_{1},\dots,\pi_{m}\rangle$ such that none of $v\circ\pi_{1},v\circ\pi_{2},\dots,v\circ\pi_{m}$ is lexicographically smaller than $v$. Such a local minimum obviously exists and can be verified efficiently (this is less obvious and we discuss it later). This means finding a local minimum is in the complexity class PLS (polynomial local search), more formally defined in the next sections. We call this problem Permutation Orbit Local Minimization, short Polm. Scheder and Tantow [SchederTantow2025] showed that Polm is PLS-complete and thus unlikely to admit an efficient algorithm. They left it as an open question whether Polm is still PLS-complete if all given permutations commute, i.e., if $\langle\pi_{1},\dots,\pi_{m}\rangle$ is Abelian. We answer this question: Yes, Abelian Polm is still PLS-complete.
@@ 1.1 Motivation
The motivation behind Polm comes from SAT solving. Kołodziejczyk and Thapen [DBLP:conf/sat/KolodziejczykT24] investigate whether proof systems including symmetry breaking rules can be simulated by proof systems without them. For a concrete example, suppose we want to prove that the Ramsey number $R(5,5)$ is at least $43$. The statement “$G$ is a graph on $43$ vertices and has no clique nor independent set of size 55” can be written as a CNF formula (here: a 10-CNF formula with $\binom{43}{2}$ variables) and fed to a SAT solver. If there is a solution, then every graph isomorphic to it is also a solution. Thus, we can try to speed things up by adding certain symmetry breaking constraints. For example, we might want to add constraints formalizing the statement “and $G$ should be lexicographically minimal”. If by “minimal” we mean “a global minimum” then this statement cannot be written efficiently as a CNF formula; if we mean “the graph cannot be made smaller by exchanging the names of two vertices” then we can encode this as a CNF formula; we are now in fact looking for a solution that is a lexicographic local minimum with respect to the permutations on the set of edges defined by swapping two vertices.
@@ 1.2 Total search problems
Another example of an FNP-problem would be Factoring, where $x$ is a number and $y$ is a prime factor of $x$, which can be checked in polynomial time. In contrast to SAT, every instance $x$ has at least one solution $y$ and hence it is a total search problem. These class of total search problems is called TFNP. If a TFNP-problem were to be NP-complete, then ${\textsc{NP}}={\textsc{coNP}}$ would follow[DBLP:journals/tcs/MegiddoP91]. Hence NP-hardness is not the right tool to study the hardness of total search problems. As TFNP is a semantic class, no complete problems are known. Instead, it is usually studied via its subclasses based on the combinatorial principle used to prove the totality of the problem. Here are some examples of subclasses, together with their defining combinatorial principle.
  • PLS: Every DAG has a sink [DBLP:journals/jcss/JohnsonPY88].
  • PPAD: If a directed graph has a vertex $u$ with more edges going out than coming in, then there must be some other vertex $v$ with more edges coming in than going out [DBLP:journals/jcss/Papadimitriou94]
  • PPA: If an undirected graph has a vertex $u$ of odd degree, it must have another such vertex $v$ [DBLP:journals/jcss/Papadimitriou94].
  • PPP: The pigeonhole principle [DBLP:journals/jcss/Papadimitriou94].
@@ 1.3 PLS
Local search is a powerful technique to solve optimization problems in practice. Nonetheless, there are instances where local search takes a long time to converge. In order to study local search from a theoretical perspective, the class PLS was introduced in 1988 [DBLP:journals/jcss/JohnsonPY88]. Following [DBLP:conf/icalp/GiannakopoulosG24], a combinatorial optimization problem can be stated as follows:
@@
PLS is now defined as the class of optimization problems where a start solution and the neighborhood function are computable in polynomial time, and the set $S$ of feasible solutions can be recognized efficiently. A local optimum is a feasible solution $s$ such that $\textnormal{cost}(s)\geq\textnormal{cost}(s^{\prime})$ for all $s^{\prime}\in N(s)$. [^2] Alternatively, we can also view it as a minimization problem, where the cost of $s$ is smaller or equal compared to all of its neighbors.
A local optimum can be found with the standard algorithm: We compute a starting solution in polynomial time and then always select an improving neighbor until no such neighbor exists. There are instances where this can take exponentially many steps[DBLP:journals/jcss/JohnsonPY88]. In fact, the standard algorithm can take exponentially long even when the optimization problem itself is in P: think about linear programming and the simplex algorithm.
PLS contains problems that are complete under search problem reductions, for example finding a locally optimal max-cut in a weighted graph[DBLP:journals/siamcomp/SchafferY91], finding locally optimal assignments for weighted CNF-formulas[DBLP:conf/coco/Krentel89] or finding pure Nash equilibria in congestion games[DBLP:conf/stoc/FabrikantPT04].
@@ 1.4 Lexicographic costs
Therefore, the costs in PLS-reductions must be large, usually of the form $2^{i}$. However, often multiple entries of the cost vector are the same. In order to compare two solutions, one would have to compute the costs of both and then compare them, which requires multiple additions and multiplications. If the cost function is like $c_{i}=2^{-i}$ and $M=2$, then we get a lexicographic cost function: $\textnormal{cost}(p_{1},x_{2})<\textnormal{cost}(p_{2},x_{2})$ if and only if $p_{1}<_{\rm lex}p_{2}$. Thus, the statement that $s=(p,x)$ is lexicographically minimal can be encoded by a simple CNF formula, which is important in contexts such as [DBLP:conf/sat/KolodziejczykT24]. [^3] For larger values of $M$ we would define lexicographic costs as $c_{i}=M^{-i}$, but we do not encounter this case in our reductions. Thus, we call a cost function lexicographic if the coefficients of the cost vector fulfill $c_{i}=C\cdot M^{-i}$, where $M$ is usually $2$, and the common factor $C$ is just for convenience, to allow for a vector $(16,8,4,2,1)$, for example. If a cost function is lexicographic, then to compare $(p,x)$ and $(p^{\prime},x^{\prime})$ it suffices to look for the first position $i$ with $p_{i}\neq p^{\prime}_{i}$; everything to the right pales in comparison. Arguably, this is one of the most simple cost function available. The central question is: Which problems are PLS-complete problems when we insist on lexicographic costs?
The first known PLS-complete problem FLIP[DBLP:journals/jcss/JohnsonPY88] fulfills this property. We are given a circuit $C$ with $n$ inputs and $m$ outputs and want to compute a locally minimal output. Formally, this optimization problem is defined by the relation $S=\{(C(x),x)\ |\ x\in\{0,1\}^{n}\}$, with the output $C(x)$ being the cost part and input $x$ being the non-cost part. The neighborhood of $(C(x),x)$ is the set of all $(C(x\oplus\mathbf{e}_{i}),x\oplus\mathbf{e}_{i})$, i.e., we are allowed to flip one bit of the input. The cost vector is $c_{i}=2^{-i}$, i.e., we want to lexicographically maximize the output. To the best of our knowledge, this has for the longest time been the only known PLS-complete problem with lexicographic costs. Recently, Polm joined this club [SchederTantow2025].
@@
Multiple PLS-complete problems are easy with a lexicographic cost function. Consider the weighted independent set problem[DBLP:journals/siamcomp/SchafferY91]. Our greedy algorithm sorts the vertices by weight from high to low and starts with $I=\emptyset$. For $i=1,\dots,n$ it adds $v_{i}$ to $I$ if $I\cup\{v_{i}\}$ is still independent. Similarly, we can find a global optimum of a max-cut instance with lexicographic weights. Here we have to check whether a given set of edges can be realized in a cut. This is easy because we just have to check whether they form a bipartite graph. More generally, Max-2-SAT with lexicographic costs can be solved efficiently because a given subset of clauses can be efficiently tested for satisfiability. It is not obvious how this works for larger clauses, as there is no known way to check satisfiability of a subset of clauses. In fact, we show that finding a local maximum under lexicographic clause weights is PLS-hard for 4-CNF formulas. We should note that even though a global lex-maximum for 2-SAT can be found efficiently, we do not know whether the standard algorithm terminates after a polynomial number of steps.
Lexicographic cost problems as starting point of PLS reductions.
Since lexicographic cost function are intuitively easier than general ones, it is not surprising that reducing from a lexicographic PLS-complete problem can be particularly easy. This is helpful if we are also interested in the hardness of approximation. As each cost entry is unique and we are only interested in the lexicographic order between two solutions, we can increase the base of the cost from 2 to be larger than the desired approximation ratio. In fact, we give a simple PLS-reduction from local max 4-SAT with lexicographic costs to computing an $\alpha$-Nash equilibrium for congestion games, thus greatly simplifying a proof by Skopalik and Vöcking [DBLP:conf/stoc/SkopalikV08].
@@ 1.5 Reductions in PLS
In order to talk about PLS-completeness, we need a notion of reductions. A PLS reduction[DBLP:journals/jcss/JohnsonPY88] from $P$ to $Q$ is a pair of functions $f$ and $g$. $f$ maps an instance $I$ of $P$ to an instance $f(I)$ of $Q$. $g$ maps a solution $s$ of $f(I)$ to a solution $g(I,s)$ of $I$, so that if $s$ is a local optimum for $f(I)$ then $g(I,s)$ is a local optimum for $I$. Additionally, both $f$ and $g$ are computable in polynomial time. One checks that this fulfills the purpose of a reduction: if we can efficiently solve $Q$ (i.e., find local minimum) then we can solve $P$. The transition graph $TG(I)$ of an instance $I$ of a PLS-problem $P$ is the directed graph where the vertices are the feasible solutions of $I$ and an edge $(u,v)$ is present if $v$ is a neighbor of $u$ according to the neighborhood function.
@@
A PLS-reduction $(f,g)$ from $P$ to $Q$ is called tight[DBLP:journals/siamcomp/SchafferY91] if for every instance $I$ of $P$ there exists a set $\mathcal{R}$ of feasible solutions of the instance $f(I)$ of $Q$ such that
@@ 1.6 Previous results
The Polm problem arose from proof-complexity based investigations into the dominance rule by Kołodziejczyk and Thapen [DBLP:conf/sat/KolodziejczykT24]. Its PLS-completeness was shown by Scheder and Tantow in [SchederTantow2025]. If the output of Polm is the locally minimized string $x\in\{0,1\}^{n}$ itself, then the problem of verifying that $x$ is in the orbit, i.e., that there is a permutation $\pi\in\langle\pi_{1},\dots,\pi_{k}\rangle$ with $x=s\circ\pi$ is known as the string isomorphism problem, which is at least as hard as graph isomorphism. Thus, to make sure Polm is in PLS, we want the output to not be $x$ but the permutation $\pi$ itself; now whether $\pi\in\langle\pi_{1},\dots,\pi_{k}\rangle$ can be verified efficiently using an algorithm of Furst, Hopcroft, and Luks [DBLP:conf/focs/FurstHL80]. For the hardness proof, they used a direct reduction from FLIP using gate gadgets and the test circuit technique by Krentel[DBLP:conf/coco/Krentel89]. While the gate gadgets could be built with commuting permutations, the test circuit technique introduced permutations that do not commute with the permutations of the gate gadgets. Hence, the question of the complexity of Polm for Abelian permutation groups arose.
Given a CNF formula with clause weights and an assignment $\alpha$, the payoff of $\alpha$ is the total weight of satisfied clauses. We denote by $k$-SAT/$d$-FLIP the optimization problem, where we are given a $k$-CNF formula with clause weights and want to find a truth assignment $\alpha$ whose payoff is locally optimal: switching $\alpha$ at at most $d$ variables does not increase the payoff. For arbitrary costs (i.e., not necessarily lexicographic) it has been known that $k$-SAT/1-FLIP is PLS-complete [DBLP:conf/focs/Krentel89]. [^4] In fact, Krentel gives a reduction to $(3,4)$-CSP and then claims that it is easy to see that it also works for SAT. Later, [DBLP:journals/siamcomp/SchafferY91] showed that Max-Cut is PLS-compete, which immediately implies that $2$-SAT/1-FLIP is PLS-complete.
Further general PLS reductions can be found in [DBLP:journals/jcss/JohnsonPY88, DBLP:journals/siamcomp/SchafferY91, DBLP:conf/coco/Krentel89, DBLP:journals/tcs/DumraufM13]. In game theory the PLS-completeness of finding a pure Nash equilibria was shown via different routes in[DBLP:conf/stoc/FabrikantPT04] and [DBLP:conf/focs/AckermannRV06]. Buchheim and Jünger proved NP-completeness for multiple problems regarding permutations in [DBLP:journals/disopt/BuchheimJ05].